# Learning Value of Information towards Joint Communication and Control in 6G V2X

### v2

@@ II System Model
We denote the headway of vehicle $i$, i.e., bumper-to-bumper distance between $i$ and its preceding vehicle (predecessor) $i-1$, by $d_{i}$, which satisfies
@@
where $L_{i-1}$ is the the length of vehicle $i-1$.
@@
At the beginning of each control interval, the control agent of vehicle $i$ has to determine the control input $u_{i,k}$ based on the observation of the state. The velocity $v_{i,k}$ and acceleration $acc_{i,k}$ can be measured locally, while the control errors $e_{pi,k}$ and $e_{vi,k}$ can be quantified by a radar unit mounted at the front of the vehicle. On the other hand, vehicle $i$ can only obtain the acceleration $acc_{i-1,k}$ of the predecessor through V2X communications. To determine the optimal action $u_{i,k}$, we must decide when and how to share information by V2X communications from predecessor $i-1$ to its follower $i$.
@@ III Preliminaries - Markov Decision Process
$v_{\pi}(s)=\mathrm{E}_{\pi}\left[\sum_{l=0}^{\infty}\gamma^{l}R_{k+l+1}\arrowvert S_{k}=s\right],\forall s\in\mathcal{S},$ (10)
@@
$q_{\pi}(s,a)=\mathrm{E}_{\pi}\left[\sum_{l=0}^{\infty}\gamma^{l}R_{k+l+1}\arrowvert S_{k}=s,a_{k}=a\right],\forall s\in\mathcal{S},a\in\mathcal{A}$ (11)
@@
$J_{\pi}=\mathrm{E}_{p(S_{0})}[v_{\pi}(S_{0})]=\mathrm{E}_{\pi,p(S_{0})}\left[\sum_{k=0}^{\infty}\gamma^{k}R_{k+1}\arrowvert S_{0}\right],$ (13)
@@ IV-A SSDP Basics
    which is derived from the predecessor $i-1$’s dynamic model $\dot{acc}_{i-1}=-\frac{1}{\rho_{i-1}}acc_{i-1}+\frac{1}{\rho_{i-1}}u_{i-1}$ based on forward Euler discretization;
@@
    The reward function reflects the optimization objective: minimizing position error $e_{pi,k}$ and velocity error $e_{vi,k}$ to achieve the platoon control target, while penalizing control input $u_{i,k}$ to reduce fuel consumption and jerk for improved driving comfort.
@@
The SSDP reduces to an MDP if and only if $\tilde{W}_{k}\subseteq\{S_{k},a_{k},\xi_{k}\}$ in Definition 1, where $\xi_{k}$ is a set of random variables whose distribution does not depend on the past states, actions and exogenous information, i.e., $\{S_{k^{\prime}},a_{k^{\prime}},W_{k^{\prime}}\}_{k^{\prime}\leq k-1}$.
@@ IV-B Augmented-State SSDP
$=f^{\tilde{S}}(\tilde{S}_{k},a_{k},\tilde{W}_{k+1}\backslash{\{\tilde{S}_{k},a_{k}\}}).$ (19)
From (2), the exogenous information becomes $\tilde{W}_{k+1}\backslash{\{\tilde{S}_{k},a_{k}\}}$, where $\tilde{W}_{k+1}\in\tilde{\mathcal{W}}$. Finally, the exogenous information function becomes $\tilde{W}_{k+1}\backslash{\{\tilde{S}_{k},a_{k}\}}=f^{\tilde{W}}(\tilde{\tilde{W}}_{k+1})$, where $\tilde{\tilde{W}}_{k+1}\in\tilde{\tilde{\mathcal{W}}}$ represents all the parameters that affect the value of $\tilde{W}_{k+1}\backslash{\{\tilde{S}_{k},a_{k}\}}$.
@@
The vehicle-following control problem relying on V2X communications where $acc_{(i-1),k}$ is transmitted from the preceding vehicle $i-1$ can be formulated as an augmented-state SSDP $\{\tilde{S}_{i,k},a_{i,k},\tilde{W}_{i,k},f^{\tilde{S}_{i}},f^{\tilde{W}_{i}},R\}$ with
@@
    where $\tilde{\tilde{W}}_{i,k}$ is the set of information that affects the control input $u_{(i-1),k}$ of predecessor $i-1$, such as the road conditions and the controller parameters of predecessor $i-1$.
@@
$=f^{\tilde{S}^{\prime}}(\tilde{S}^{\prime}_{k},a_{k},\tilde{\tilde{W}}_{k+1}\backslash\{\tilde{S}^{\prime}_{k},a_{k}\}).$ (22)
From (3), the exogenous information becomes $\tilde{\tilde{W}}_{k+1}\backslash\{\tilde{S}^{\prime}_{k},a_{k}\}$, where $\tilde{\tilde{W}}_{k+1}\in\tilde{\tilde{\mathcal{W}}}$. Finally, the exogenous information function becomes $\tilde{\tilde{W}}_{k+1}\backslash\{\tilde{S}^{\prime}_{k},a_{k}\}=f^{\tilde{\tilde{W}}}(\tilde{\tilde{\tilde{W}}}_{k+1})$, where $\tilde{\tilde{\tilde{W}}}_{k+1}\in\tilde{\tilde{\tilde{\mathcal{W}}}}$ represents all the parameters that affect the value of $\tilde{\tilde{W}}_{k+1}\backslash\{\tilde{S}^{\prime}_{k},a_{k}\}$.
@@ V-A Utility-based VoI
(Value of Imperfect Information (VoII)) The VoII is defined as the difference in performance between the acting policy $\tilde{\pi}^{\rm ip}$ for the augmented-state SSDP with imperfect information, where the state $\tilde{S}_{k}^{\mathrm{ip}}$ is constructed from imperfect information $I_{k}^{\rm ip}$, and the optimal policy $\tilde{\pi}^{*}$ for the augmented-state SSDP $\tilde{S}_{k}=\{S_{k},I_{k}\}$, where the augmented state $\tilde{S}_{k}=\{S_{k},I_{k}\}$ includes perfect information $I_{k}$.
@@
[…]
@@
Therefore, utility-based VoI can be categorized into two main types: EVoI and IVoI. Each of these types is further divided into two subcategories: VoMI and VoII. As illustrated in Fig. 3, this classification results in four distinct categories of utility-based VoI: Expected VoMI (EVoMI), Expected VoII (EVoII), Immediate VoMI (IVoMI), and Immediate VoII (IVoII). According to Definition 5 and 6, the policy/state/action $\pi^{\rm{inf}}$/$S_{k}^{\rm{inf}}$/$a_{k}^{\rm{inf}}$ in (23) and (24) is replaced by $\pi$/$S_{k}$/$a_{k}$ and $\tilde{\pi}^{\rm ip}$/$\tilde{S}_{k}^{\rm ip}$/$a_{k}^{\rm ip}$ for VoMI and VoII, respectively, while $\pi^{\rm{sup}}$/$S_{k}^{\rm{sup}}$ in (23) and (24) is replaced by $\tilde{\pi}^{*}$/$\tilde{S}_{k}$.
@@
Consider the vehicle-following control problem in Case 1. The control agent of vehicle $i$ can observe its position error $e_{pi,k}$, velocity error $e_{vi,k}$, and acceleration $acc_{i,k}$ based on local sensory information. Meanwhile, the acceleration $acc_{i-1,k}$ of the predecessor $i-1$ is remote sensory information that has to be transmitted from vehicle $i-1$ to vehicle $i$. In order to quantify the benefit of making decisions when $acc_{i-1,k}$ is available to the control agent of vehicle $i$, the four utility-based VoI can be defined for $acc_{i-1,k}$:
@@
  • EVoMI $\Xi_{i}^{\rm M}=J_{\pi_{i}}-J_{\tilde{\pi}_{i}^{*}}$, where $J_{\pi_{i}}$ and $J_{\tilde{\pi}_{i}^{*}}$ are the performance of the acting policy $\pi_{i}$ based on partial observations $S_{i,k}=[e_{pi,k},e_{vi,k},acc_{i,k}]^{\mathrm{T}}$ (for the SSDP in Case 2) and the optimal policy $\tilde{\pi}_{i}^{*}$ based on full observations $\tilde{S}_{i,k}=[e_{pi,k},e_{vi,k},acc_{i,k},acc_{(i-1),k}]^{\mathrm{T}}$ (for the augmented-state SSDP in Case 4), respectively. The EVoMI quantifies the performance loss incurred when the following vehicle $i$ consistently fails to receive the acceleration $acc_{i-1,k}$ of the preceding vehicle $i-1$ at every control interval $k$ throughout the entire episode. Note that the performance refers to the expected cumulative reward defined in (13), while the reward function of the vehicle-following control problem, given in (2), depends on the position error, velocity error, and acceleration of vehicle $i$.
@@
  • IVoMI $\xi_{i,k}^{\rm M}=A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})|_{u_{i,k}\sim\pi_{i}(\cdot|S_{i,k})}$, where $A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})$ is the advantage function of the optimal policy $\tilde{\pi}_{i}^{*}$, and $u_{i,k}\sim\pi_{i}(\cdot|S_{i,k})$ is the control input of vehicle $i$ that is sampled according to the acting policy $\pi_{i}(\cdot|S_{i,k})$ based on partial observation $S_{i,k}$. The IVoMI quantifies the performance loss incurred when the following vehicle $i$ fails to receive the acceleration $acc_{i-1,k}$ of the preceding vehicle $i-1$ at a specific control interval $k$, while consistently receiving this information during the rest of the episode. Consequently, the IVoMI is a function of the state action pair $\{\tilde{S}_{i,k},u_{i,k}\}$, where $\tilde{S}_{i,k}$ represents the full observation at control interval $k$, including $acc_{i-1,k}$. It measures the performance loss incurred when the control agent has access only to the partial observation (excluding $acc_{i-1,k}$) and selects action $u_{i,k}$ using the acting control policy. Unlike EVoI, IVoI distinguishes the value of knowing the predecessor’s acceleration based on both its specific values and the available system state information, such as position error, velocity error, and acceleration of the follower.
@@
  • EVoII $\Xi_{i}^{\rm I}=J_{\pi_{i}^{\rm d}}-J_{\tilde{\pi}_{i}^{*}}$, where $J_{\pi_{i}^{\rm d}}$ and $J_{\tilde{\pi}_{i}^{*}}$ are the performance of the acting policy $\pi_{i}^{\rm d}$ based on imperfect (delayed) observations $\tilde{S}_{i,k}^{\rm d}=[x_{i,k-\tau_{i,k}},acc_{(i-1),k-\tau_{i,k}},\{u_{i,k^{\prime}}\}_{k^{\prime}=k-\tau_{\mathrm{max}}}^{k-1},\tau_{i,k}]^{\mathrm{T}}$ (for the Random Delay SSDP in Case 5) and the optimal policy $\tilde{\pi}_{i}^{*}$ based on perfect observations $\tilde{S}_{i,k}=[e_{pi,k},e_{vi,k},acc_{i,k},acc_{(i-1),k}]^{\mathrm{T}}$ (for the augmented-state SSDP in Case 4), respectively. The EVoII is similar to the EVoMI, with the only difference being that the performance loss results from the following vehicle $i$ receiving a delayed acceleration $acc_{i-1,k-\tau_{k}}$ from the predecessor $i-1$, rather than not receiving it at all.
@@ V-B Information-Theory-based VoI
Note that the KL divergence in (9) and (9) is a measure of the information lost when $T^{S}\otimes T^{I}$ and $T^{r(S)}$ are used for approximating $T^{\tilde{S}}$ and $T^{r(\tilde{S})}$, respectively. The ITVoI is 00 if the transition of $S_{k}$ and the reward $R_{k+1}$ are independent of $I_{k}$. In this case, we know from Lemma 1 that the performance of the optimal policies of the augmented-state SSDP and the original SSDP are the same, and there is no need to include $I_{k}$ in $\tilde{S}_{k}$. On the other hand, a higher KL divergence value indicates that the transition of $S_{k}$ or the reward $R_{k+1}$ depend on $I_{k}$ to a larger extent, and thus the availability of $I_{k}$ is more important for accurately predicting the future state $S_{k+1}$ or predicting the reward $R_{k+1}$. In this case, including $I_{k}$ in $\tilde{S}_{k}$ will be more likely to improve the performance of the optimal policy. Therefore, the KL divergence is a suitable quantitative measure for the VoI.
@@
From (7), we can see that the ITVoI of $acc_{i-1,k}$ depends on the conditional probability $p\{acc_{i-1,k}|S_{i,k},u_{i,k}\}$, i.e., how well we can predict the acceleration $acc_{i-1,k}$ of predecessor $i-1$ given the driving status $S_{i,k}$ and control input $u_{i,k}$ of the following vehicle $i$? Consider the extreme case when $acc_{i-1,k}=g(S_{i,k},u_{i,k})$ is a deterministic function of $S_{i,k}$ and/or $u_{i,k}$, we thus have $p\{acc_{i-1,k}|S_{i,k},u_{i,k}\}=\mathbf{1}_{acc_{i-1,k}=g(S_{i,k},u_{i,k})}$. Here, the ITVoI of $acc_{i-1,k}$ is 00, meaning there is no value in knowing $acc_{i-1,k}$. Note that this scenario is not possible in practice, since $acc_{i-1,k}$ is a function of $acc_{i-1,k-1}$ and $u_{i-1,k-1}$ according to (2). The predecessor $i-1$ cannot determine its control input $u_{i-1,k-1}$ at control interval $k-1$ based on $S_{i,k}$ and/or $u_{i,k}$, whose values are not available until control interval $k$. In the other extreme case, if the predecessor $i-1$ makes driving decisions completely independently of the following vehicle $i$, we have $p\{acc_{i-1,k}|S_{i,k},u_{i,k}\}=p\{acc_{i-1,k}\}$. This means the ITVoI depends on the probability distribution of $acc_{i-1,k}$. For example, if $acc_{i-1,k}$ is a constant value, i.e., $p\{acc_{i-1,k}=c\}=1$, the ITVoI is 00.

### v3

@@ II System Model
We denote the headway of vehicle $i$, i.e., bumper-to-bumper distance between $i$ and its preceding vehicle (predecessor) i1i-1, by $d_{i}$, which satisfies
@@
where $L_{i-1}$ is the the length of vehicle i1i-1.
@@
At the beginning of each control interval, the control agent of vehicle $i$ has to determine the control input $u_{i,k}$ based on the observation of the state. The velocity $v_{i,k}$ and acceleration $acc_{i,k}$ can be measured locally, while the control errors $e_{pi,k}$ and $e_{vi,k}$ can be quantified by a radar unit mounted at the front of the vehicle. On the other hand, vehicle $i$ can only obtain the acceleration $acc_{i-1,k}$ of the predecessor through V2X communications. To determine the optimal action $u_{i,k}$, we must decide when and how to share information by V2X communications from predecessor i1i-1 to its follower $i$.
@@ III Preliminaries - Markov Decision Process
[…]
@@
[…]
@@
[…]
@@ IV-A SSDP Basics
    which is derived from the predecessor i1i-1’s dynamic model $\dot{acc}_{i-1}=-\frac{1}{\rho_{i-1}}acc_{i-1}+\frac{1}{\rho_{i-1}}u_{i-1}$ based on forward Euler discretization;
@@
    The reward function reflects the optimization objective: minimizing position error $e_{pi,k}$ and velocity error $e_{vi,k}$ to achieve the vehicle following target, while penalizing control input $u_{i,k}$ to reduce fuel consumption and jerk for improved driving comfort.
@@
The SSDP reduces to an MDP if and only if $\tilde{W}_{k}\subseteq\{S_{k},a_{k},\zeta_{k}\}$ in Definition 1, where $\zeta_{k}$ is a set of random variables whose distribution does not depend on the past states, actions and exogenous information, i.e., $\{S_{k^{\prime}},a_{k^{\prime}},W_{k^{\prime}}\}_{k^{\prime}\leq k-1}$.
@@ IV-B Augmented-State SSDP
[…]
@@
The vehicle-following control problem relying on V2X communications where $acc_{(i-1),k}$ is transmitted from the preceding vehicle i1i-1 can be formulated as an augmented-state SSDP $\{\tilde{S}_{i,k},a_{i,k},\tilde{W}_{i,k},f^{\tilde{S}_{i}},f^{\tilde{W}_{i}},R\}$ with
@@
    where $\tilde{\tilde{W}}_{i,k}$ is the set of information that affects the control input $u_{(i-1),k}$ of predecessor i1i-1, such as the road conditions and the controller parameters of predecessor i1i-1.
@@
[…]
@@ V-A Utility-based VoI
(Value of Imperfect Information (VoII)) The VoII is defined as the difference in performance between the acting policy $\tilde{\pi}^{\rm ip}$ for the augmented-state SSDP with imperfect information, where the state $\tilde{S}_{k}^{\mathrm{ip}}$ is constructed from imperfect information $I_{k}^{\rm ip}$, and the optimal policy $\tilde{\pi}^{*}$ for the augmented-state SSDP, where the augmented state $\tilde{S}_{k}=\{S_{k},I_{k}\}$ includes perfect information $I_{k}$.
@@
$\bar{\Xi}=\mathrm{E}_{\pi^{\rm CM}}\mathrm{E}_{\pi^{\rm inf}}\left[\sum_{k=0}^{\infty}\gamma^{k}\xi_{k}\right],\ 0\leq\gamma\leq 1,$ (25)
@@
Therefore, utility-based VoI can be categorized into two main types: EVoI and IVoI. Each of these types is further divided into two subcategories: VoMI and VoII. As illustrated in Fig. 3, this classification results in four distinct categories of utility-based VoI: Expected VoMI (EVoMI), Expected VoII (EVoII), Immediate VoMI (IVoMI), and Immediate VoII (IVoII). According to Definition 5 and 6, the policy/state/action $\pi^{\rm{inf}}$$S_{k}^{\rm{inf}}$$a_{k}^{\rm{inf}}$ in (23) and (24) is replaced by $\pi$$S_{k}$$a_{k}$ and $\tilde{\pi}^{\rm ip}$$\tilde{S}_{k}^{\rm ip}$$a_{k}^{\rm ip}$ for VoMI and VoII, respectively, while $\pi^{\rm{sup}}$$S_{k}^{\rm{sup}}$ in (23) and (24) is replaced by $\tilde{\pi}^{*}$$\tilde{S}_{k}$.
@@
Consider the vehicle-following control problem in Case 1. The control agent of vehicle $i$ can observe its position error $e_{pi,k}$, velocity error $e_{vi,k}$, and acceleration $acc_{i,k}$ based on local sensory information. Meanwhile, the acceleration $acc_{i-1,k}$ of the predecessor i1i-1 is remote sensory information that has to be transmitted from vehicle i1i-1 to vehicle $i$. In order to quantify the benefit of making decisions when $acc_{i-1,k}$ is available to the control agent of vehicle $i$, the four utility-based VoI can be defined for $acc_{i-1,k}$:
@@
  • EVoMI $\Xi_{i}^{\rm M}=J_{\pi_{i}}-J_{\tilde{\pi}_{i}^{*}}$, where $J_{\pi_{i}}$ and $J_{\tilde{\pi}_{i}^{*}}$ are the performance of the acting policy $\pi_{i}$ based on partial observations $S_{i,k}=[e_{pi,k},e_{vi,k},acc_{i,k}]^{\mathrm{T}}$ (for the SSDP in Case 2) and the optimal policy $\tilde{\pi}_{i}^{*}$ based on full observations $\tilde{S}_{i,k}=[e_{pi,k},e_{vi,k},acc_{i,k},acc_{(i-1),k}]^{\mathrm{T}}$ (for the augmented-state SSDP in Case 4), respectively. The EVoMI quantifies the performance loss incurred when the following vehicle $i$ consistently fails to receive the acceleration $acc_{i-1,k}$ of the preceding vehicle i1i-1 at every control interval $k$ throughout the entire episode. Note that the performance refers to the expected cumulative reward defined in (13), while the reward function of the vehicle-following control problem, given in (2), depends on the position error, velocity error, and acceleration of vehicle $i$.
@@
  • IVoMI $\xi_{i,k}^{\rm M}=A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})|_{u_{i,k}\sim\pi_{i}(\cdot|S_{i,k})}$, where $A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})$ is the advantage function of the optimal policy $\tilde{\pi}_{i}^{*}$, and $u_{i,k}\sim\pi_{i}(\cdot|S_{i,k})$ is the control input of vehicle $i$ that is sampled according to the acting policy $\pi_{i}(\cdot|S_{i,k})$ based on partial observation $S_{i,k}$. The IVoMI quantifies the performance loss incurred when the following vehicle $i$ fails to receive the acceleration $acc_{i-1,k}$ of the preceding vehicle i1i-1 at a specific control interval $k$, while consistently receiving this information during the rest of the episode. Consequently, the IVoMI is a function of the state action pair $\{\tilde{S}_{i,k},u_{i,k}\}$, where $\tilde{S}_{i,k}$ represents the full observation at control interval $k$, including $acc_{i-1,k}$. It measures the performance loss incurred when the control agent has access only to the partial observation (excluding $acc_{i-1,k}$) and selects action $u_{i,k}$ using the acting control policy. Unlike EVoI, IVoI distinguishes the value of knowing the predecessor’s acceleration based on both its specific values and the available system state information, such as position error, velocity error, and acceleration of the follower.
@@
  • EVoII $\Xi_{i}^{\rm I}=J_{\pi_{i}^{\rm d}}-J_{\tilde{\pi}_{i}^{*}}$, where $J_{\pi_{i}^{\rm d}}$ and $J_{\tilde{\pi}_{i}^{*}}$ are the performance of the acting policy $\pi_{i}^{\rm d}$ based on imperfect (delayed) observations $\tilde{S}_{i,k}^{\rm d}=[x_{i,k-\tau_{i,k}},acc_{(i-1),k-\tau_{i,k}},\{u_{i,k^{\prime}}\}_{k^{\prime}=k-\tau_{\mathrm{max}}}^{k-1},\tau_{i,k}]^{\mathrm{T}}$ (for the Random Delay SSDP in Case 5) and the optimal policy $\tilde{\pi}_{i}^{*}$ based on perfect observations $\tilde{S}_{i,k}=[e_{pi,k},e_{vi,k},acc_{i,k},acc_{(i-1),k}]^{\mathrm{T}}$ (for the augmented-state SSDP in Case 4), respectively. The EVoII is similar to the EVoMI, with the only difference being that the performance loss results from the following vehicle $i$ receiving a delayed acceleration $acc_{i-1,k-\tau_{k}}$ from the predecessor i1i-1, rather than not receiving it at all.
@@ V-B Information-Theory-based VoI
Note that the KL divergence in (9) and (9) is a measure of the information lost when $T^{S}\otimes T^{I}$ and $T^{r(S)}$ are used for approximating $T^{\tilde{S}}$ and $T^{r(\tilde{S})}$, respectively. The ITVoI is 0 if the transition of $S_{k}$ and the reward $R_{k+1}$ are independent of $I_{k}$. In this case, we know from Lemma 1 that the performance of the optimal policies of the augmented-state SSDP and the original SSDP are the same, and there is no need to include $I_{k}$ in $\tilde{S}_{k}$. On the other hand, a higher KL divergence value indicates that the transition of $S_{k}$ or the reward $R_{k+1}$ depend on $I_{k}$ to a larger extent, and thus the availability of $I_{k}$ is more important for accurately predicting the future state $S_{k+1}$ or predicting the reward $R_{k+1}$. In this case, including $I_{k}$ in $\tilde{S}_{k}$ will be more likely to improve the performance of the optimal policy. Therefore, the KL divergence is a suitable quantitative measure for the VoI.
@@
From (7), we can see that the ITVoI of $acc_{i-1,k}$ depends on the conditional probability $p\{acc_{i-1,k}|S_{i,k},u_{i,k}\}$, i.e., how well we can predict the acceleration $acc_{i-1,k}$ of predecessor i1i-1 given the driving status $S_{i,k}$ and control input $u_{i,k}$ of the following vehicle $i$? Consider the extreme case when $acc_{i-1,k}=g(S_{i,k},u_{i,k})$ is a deterministic function of $S_{i,k}$ and/or $u_{i,k}$, we thus have $p\{acc_{i-1,k}|S_{i,k},u_{i,k}\}=\mathbf{1}_{acc_{i-1,k}=g(S_{i,k},u_{i,k})}$. Here, the ITVoI of $acc_{i-1,k}$ is 0, meaning there is no value in knowing $acc_{i-1,k}$. Note that this scenario is not possible in practice, since $acc_{i-1,k}$ is a function of $acc_{i-1,k-1}$ and $u_{i-1,k-1}$ according to (2). The predecessor i1i-1 cannot determine its control input $u_{i-1,k-1}$ at control interval k1k-1 based on $S_{i,k}$ and/or $u_{i,k}$, whose values are not available until control interval $k$. In the other extreme case, if the predecessor i1i-1 makes driving decisions completely independently of the following vehicle $i$, we have $p\{acc_{i-1,k}|S_{i,k},u_{i,k}\}=p\{acc_{i-1,k}\}$. This means the ITVoI depends on the probability distribution of $acc_{i-1,k}$. For example, if $acc_{i-1,k}$ is a constant value, i.e., $p\{acc_{i-1,k}=c\}=1$, the ITVoI is 0.